Southern Rim Parts estimates its manufacturing overhead to be $318,000 and its direct labor costs to be $1,060,000 for year 1. The first three jobs that Southern Rim worked on had actual direct labor costs of $65,000 for Job 301, $90,000 for Job 302, and $175,000 for Job 303. For the year, actual manufacturing overhead was $399,000 and total direct labor cost was $834,000. Manufacturing overhead is applied to jobs on the basis of direct labor costs using predetermined rates.



Required:


a. How much overhead was assigned to each of the three jobs, 301, 302, and 30?


 


b. What was the over- or underapplied manufacturing overhead for year 1?
Answer #1

a)

Overhead application rate = Budgeted Overhead / Application Base

= Budgeted Overhead / Budgeted Direct Labor costs

= 318,000/1,060,000

= 0.3 per dollar of direct labor cost.

Overhead assignment to jobs = Budgeted rate x Labor Cost.

Job 301

= 0.3 * 65,000 = 19,500

Job 302

= 0.3 * 90,000 = 27,000

Job 303

= 0.3 * 175,000 = 52,500

b)

Overhead applied for the year = Total direct labor cost x Overhead rate

= 834,000*0.3

= 250,200

Under-applied overhead = Actual overhead - Applied overhead

= 399,000-250,200

= 148,800